# Gradient Inversion Attack on Graph Neural Networks

### v1

@@ Abstract
Graph federated learning is of essential importance for training over large graph datasets while protecting data privacy, where each client stores a subset of local graph data, while the server collects the local gradients and broadcasts only the aggregated gradients. Recent studies reveal that a malicious attacker can steal private image data from gradient exchanging of neural networks during federated learning. However, none of the existing works have studied the vulnerability of graph data and graph neural networks under such attack. To answer this question, the present paper studies the problem of whether private data can be recovered from leaked gradients in both node classification and graph classification tasks and proposes a novel attack named Graph Leakage from Gradients (GLG). Two widely-used GNN frameworks are analyzed, namely GCN and GraphSAGE. The effects of different model settings on recovery are extensively discussed. Through theoretical analysis and empirical validation, it is shown that parts of the graph data can be leaked from the gradients.
@@ 1 Introduction
Federated Learning McMahan et al. (2016b) is a distributed learning paradigm that has gained increasing attention. Various organizations in healthcare rely on federated learning to train models while also keeping user data private. Gradient averaging is a widely used mechanism in federated learning, where clients send the gradients of the models instead of the actual private data to a central server. The server then updates the model by taking the average gradient over all the clients. The computation is executed independently on each client and synchronized via exchanging gradients between the server Li et al. (2014); Iandola et al. (2016) and the clients Patarasuk & Yuan (2009). It enables the training of a global model using data distributed at multiple sources without the need to send them to a central location. This scheme is widely used when the training set is large and/or contains private information  Jochems et al. (2016); Konečný et al. (2016). For example, multiple hospitals train a model jointly without sharing their patients’ medical records Jochems et al. (2017); McMahan et al. (2016a).
Recently, several efforts have demonstrated that gradients can leak input data through gradient inversion attacks.  Jeon et al. (2021); Zhu et al. (2019); Geiping et al. (2020); Zhao et al. (2020b); Yin et al. (2021); Zhu & Blaschko (2020); Jin et al. (2021) successfully reconstruct the users’ private image data leveraging the gradients. Specifically, Zhao et al. (2020a) proposes that the ground truth labels can be directly inferred from the gradients. Consequently, ensuing works improve the attack techniques on large batches of user data Yin et al. (2021); Fowl et al. (2022b). A few recent works study recovering text data in federated learning. To facilitate the text recovery,  Boenisch et al. (2021); Fowl et al. (2022a) recover text data using a strong threat model in which the server is malicious and is able to manipulate the training model’s weights, while Gupta et al. (2022) studies the feasibility of recovering text from large batch sizes by leveraging the memorization capability of Language Models during federated learning. It can be seen that the majority of the works are based on either image or text data.
@@
On another note, Graph Neural Networks (GNNs) have demonstrated state-of-the-art results when applied to structured data. From learning meaningful node representations over large social networks Hamilton et al. (2017), to predicting molecular property Gómez-Bombarelli et al. (2018) Yang et al. (2019), GNNs have been widely adopted across various paradigms. The confluence of Federated Learning and graphs has led to the emergence of Federated Graph Learning He et al. (2021) Scardapane et al. (2020) Zhang et al. (2021). Despite the notable success of gradient inversion attacks for general deep neural networks, none of existing works studied the vulnerability of GNNs towards gradient inversion attacks. Hence, the present work aims to answer the following unexplored question:
Can graph data be leaked from gradients of GNNs?
@@
A major challenge in recovering graph data is the intertwined nature of the nodal features and the irregular graph structures, making it difficult to separate the information from the two sources. In other words, representations are correlated, and the underlying graph structure influences the correlation. In this work, we first provide a theoretical analysis of the feasibility of recovering graph data based on gradients. Motivated by this observation, we propose a gradient inversion attack called GLG (Graph Leakage From Gradients). We investigate two commonly used GNN frameworks, i.e., Graph Convolutional Networks (GCN) Kipf & Welling (2016), and GraphSAGE Hamilton et al. (2017), and examine their vulnerabilities to gradient-based attacks for both node and graph classification tasks. An illustration of the attack scenario is presented in Figure 1, showcasing the process in which a curious central parameter server attempts to reconstruct the graph data utilizing the gradients of the model obtained from a client. To our best knowledge, this is the first work that studies gradient-based attacks for graph data. Our main contributions are summarized as follows:
@@
i) Theoretical analysis highlighting the vulnerabilities of the GNN frameworks towards gradient-based attack.
@@
ii) Developing novel gradient inversion attack mechanisms to recover graph data for both node and graph classification tasks.
@@
iii) Experiments are conducted in a variety of settings with different amounts of prior knowledge.
@@ 2.1 GNN Frameworks
To pave the road for ensuing analysis and algorithm design, we first introduce GraphSAGE and Graph Convolutional Network (GCN) frameworks for both node and graph classification tasks below.
@@ 3 Gradient Inversion Attack for GNNs
We consider the scenario where there is a curious but honest server, i.e., the server has access to the model gradient with respect to the client’s private graph data. Different attack scenarios are extensively discussed where the server may also have access to different amounts of information.
@@ 3.1 Threat Model
The attacker is an honest-but-curious server that aims to recover the user’s graph data given the gradients of a GNN model. We thoroughly investigate attacks targeting both node classification and graph classification tasks. In addition to possessing knowledge of the gradients, the attacker may also have access to varying amounts of information about the user’s private graph data. Hence we categorize the attacks based on the task as follows:
Node Attacker 1: has access to the gradients of a node classification model for the target node and recovers the nodal features of the target node and its neighbors.
Node Attacker 2: has access to the gradients of a node classification model for all the nodes in an egonet/subgraph. Specifically, the gradients are obtained for the loss function $\mathcal{L}_{v}\in\mathbb{R}^{1\times N}$ with respect to the neural networks weights, where each element $\mathcal{L}_{v_{i}}$ denotes the loss for node $v_{i}$. Unlike traditional classification tasks, the adjacency information is also significant for graph data. In this setting, the graph structure could be unknown. The honest-but-curious server would also try to recover the nodal features and/or the graph structure. Based on the input, we categorize the attacker in more detail as follows
  • Node Attacker 2-a: has access to the gradients of all the nodes in an egonet/subgraph along with the nodal features and recovers the underlying graph structure.
  • Node Attacker 2-b: has access to the gradients of all the nodes in an egonet/subgraph along with the graph structure and recovers the nodal features.
  • Node Attacker 2-c: has access only to the gradients of all nodes in an egonet/subgraph and recovers the graph structure as well as the nodal features.
Graph Attacker: has access to the gradients of the graph for a graph classification model. Similar to Node Attacker 2, in this setting the honest-but-curious server would try to recover the nodal features and/or the graph structure. Hence, we categorize the attack in more detail as follows
  • Graph Attacker-a: has access to the gradients along with the nodal features and recovers the underlying graph structure .
  • Graph Attacker-b: has access to the gradients along with the graph structure and recovers the nodal features.
  • Graph Attacker-c: has access to the gradients and recovers both the nodal features and the graph structure.
@@ 3.2 Attack Mechanisms
In this section, GLG is introduced to attack a GNN framework and steal private user data using the leaked gradients in a federated graph learning setting. Specifically, given the gradients of the model from a client for the $i$-th data sample $\nabla_{\mathbf{W}}\mathcal{L}_{c,i}$, the goal is to recover the input graph data sample that was fed to the model. For simplicity, we denote $\nabla_{\mathbf{W}}\mathcal{L}_{c,i}$ as $\nabla_{\mathbf{W}}\mathcal{L}$ for the graph classification task, and as $\nabla_{\mathbf{W}}\mathcal{L}_{v}$ for the node classification task. In Zhu et al. (2019), the input data to a model is recovered by matching the dummy gradients with the leaked gradients. The dummy gradients are obtained by feeding dummy input data to the model. Given the gradients at a certain time step as $\nabla_{\mathbf{W}}\mathcal{L}$, the input data is recovered by minimizing the following objective function
@@
In this work, it is assumed that the input label is known since in a classification task with cross-entropy loss, the labels can be readily inferred from the gradients Zhao et al. (2020a). For further details regarding this please refer to Appendix B. Many real-world graphs, such as social networks have the property that connected nodes are similar to each other (feature smoothness). To ensure the smoothness of the reconstructed graph data, the following loss function can be used similar to Zhang et al. (2022).
$\mathcal{L}_{s}$ $=tr(\mathbf{X}\mathbf{L}\mathbf{X}^{\top})$
$=\sum_{i,j\in\mathcal{E}}(\frac{\mathbf{x}_{v_{i}}}{\sqrt{d_{i}}}-\frac{\mathbf{x}_{v_{j}}}{\sqrt{d_{j}}})$ (13)
@@
where $\mathbb{D}$ is defined in equation 12, $\alpha$ and $\beta$ are hyperparameters. The iterative algorithm for recovering the private data from the gradients in a node classification task and a graph classification task are summarized in algorithms 1, 2 and 3, respectively. Specifically, Node Attacker 1 utilizes algorithm 1 for recovering the nodal features of the target node and its neighbors. Node Attackers 2 (a, b and c) utilizes algorithm 2 for recovering the nodal features and graph structure given the gradients of the loss function for each node in the egonet/subgraph. By default, algorithm 2 assumes that both nodal features and graph structure are unknown and tries to recover both. However, only the unknown parameter will be optimized when either of the two is known. For instance, Node Attacker 2-a only optimizes for $\mathbf{A}$ in Line 9 of algorithm 2 while skipping Line 8 since the nodal features $\mathbf{X}$ is known. The same holds for Node Attacker 2-b. However, unlike the other two Node Attacker 2-c optimizes for both. Similar logic holds for Graph Attacker (a, b and c) which utilizes algorithm 3. Note that algorithms 2 and 3 might seem similar since they are recovering the same variables. However, a key difference between the two is that algorithm 2 is attacking a node classification model while algorithm 3 attacks a graph classification model. In algorithm 2, the gradients of the loss function are computed individually for each node. This is because each node within the subgraph has an associated label. In contrast, algorithm 3 calculates gradients for the entire graph, as the graph has a single label.
Recovering the Adjacency Matrix: In all the scenarios where the attacker is trying to recover the adjacency matrix, projected gradient descent is applied. Specifically, the gradient descent step in Line 9 of algorithms 2 and 3 entails a projection step, where each entry $(i,j)$ of the adjacency matrix is updated through the entry-wise projection operator defined as
@@
Finally, to recover the binary adjacency matrix, we consider each entry $\hat{\mathbf{A}}_{ij}$ as the probability of any edge between nodes $v_{i}$ and $v_{j}$ and the recovered binary adjacency matrix is obtained by sampling from a Bernoulli distribution with the corresponding probability at the last iteration, see also Line 11 of algorithm 2 and algorithm 3.
Since the recovered adjacency matrix $\hat{\mathbf{A}}$ is obtained via sampling through Bernoulli distribution, the Frobenius regularizer has the effect of reducing the magnitude of entries of $\hat{\mathbf{A}}$. Consequently, this regularizer also contributes to the promotion of sparsity in the ultimately recovered adjacency matrix. This sparsity property is particularly relevant since real-world social networks are sparse.

### v2

@@ Abstract
Graph federated learning is of essential importance for training over large graph datasets while protecting data privacy, where each client stores a subset of local graph data, while the server collects the local gradients and broadcasts only the aggregated gradients. Recent studies reveal that a malicious attacker can steal private image data from the gradient exchange of neural networks during federated learning. However, the vulnerability of graph data and graph neural networks under such attacks, i.e., reconstructing both node features and graph structure from gradients, remains largely underexplored. To answer this question, this paper studies the problem of whether private data can be reconstructed from leaked gradients in both node classification and graph classification tasks and proposes a novel attack named Graph Leakage from Gradients (GLG). Two widely used GNN frameworks are analyzed, namely GCN and GraphSAGE. The effects of different model settings on reconstruction are extensively discussed. Theoretical analysis and empirical validation demonstrate that, by leveraging the unique properties of graph data and GNNs, GLG achieves more accurate reconstruction of both nodal features and graph structure from gradients.
@@ 1 Introduction
Federated Learning (McMahan et al., 2016b) is a distributed learning paradigm that has gained increasing attention. Gradient averaging is a widely used mechanism in federated learning, where clients send the gradients of the models instead of the actual private data to a central server. The server then updates the model by taking the average gradient over all the clients. The computation is executed independently on each client and synchronized via exchanging gradients between the server (Li et al., 2014; Iandola et al., 2016) and the clients (Patarasuk & Yuan, 2009). Various organizations and mobile devices in healthcare rely on federated learning to train models while also keeping user data private. For example, multiple hospitals train a model jointly without sharing their patients’ medical records (Jochems et al., 2017; McMahan et al., 2016a).
Recently, several efforts have demonstrated that gradients, accessible to an honest-but-curious (HBC) server, can be exploited via gradient inversion attacks to reconstruct the original input data. By leveraging the gradients, Jeon et al. (2021); Zhu et al. (2019); Geiping et al. (2020); Zhao et al. (2020b); Yin et al. (2021); Zhu & Blaschko (2020); Jin et al. (2021) successfully reconstruct the users’ private image data and Boenisch et al. (2021); Fowl et al. (2022a) reconstruct users’ text data. Gradient inversion attacks have demonstrated the privacy vulnerabilities of standard federated learning.
@@
Meanwhile, Graph Neural Networks (GNNs) have demonstrated state-of-the-art results when applied to structured data. From learning meaningful node representations over large social networks (Hamilton et al., 2017), to predicting molecular property (Gómez-Bombarelli et al., 2018; Yang et al., 2019), GNNs have been widely adopted across various paradigms. The confluence of Federated Learning and graphs has led to the emergence of Federated Graph Learning (FGL) (He et al., 2021; Scardapane et al., 2020; Zhang et al., 2021).
However, the privacy concerns regarding gradient leakage in FGL remain crucial and underexplored. Prior work has explored gradient-based attacks and privacy techniques in federated graph learning (Geisler et al., 2021; Ge et al., 2024), which aims to mislead training. Meanwhile, gradient inversion attacks have shown notable success in general deep neural networks (Zhu et al., 2019; Zhao et al., 2020a). Nevertheless, graph-specific gradient-inversion attacks, i.e., reconstructing both node features and graph structure from gradients, remain largely underexplored. Hence, the present work aims to answer the following underexplored research question:
How can the unique properties of graph data and GNNs be leveraged to effectively reconstruct graph structure and node features from gradients of GNNs?
@@
A major challenge in reconstructing graph data is the intertwined nature of the nodal features and the irregular graph structures, making it difficult to separate the information from nodal features and the graph structure. In other words, representations are correlated, and the underlying graph structure influences the correlation. In this work, we provide a theoretical analysis of the feasibility of reconstructing graph data based on gradients. Motivated by this observation, we propose a gradient inversion attack called GLG (Graph Leakage From Gradients). We investigate two commonly used GNN frameworks, i.e., Graph Convolutional Networks (GCN) (Kipf & Welling, 2016), and GraphSAGE (Hamilton et al., 2017), and examine their vulnerabilities to gradient inversion attacks for both node and graph classification tasks. An illustration of the attack scenario is presented in Figure 1, showcasing the process in which a curious central parameter server attempts to reconstruct the graph data utilizing the gradients of the model obtained from a client. To our best knowledge, this is among the first works to study gradient-inversion attacks for graph data. Our main contributions are summarized as follows:
@@
  • Theoretical analysis revealing the vulnerabilities of GNNs towards gradient inversion attack.
@@
  • Developing novel gradient inversion attack mechanisms to reconstruct graph data for both node and graph classification tasks.
@@
  • Experiments are conducted in a variety of settings with different amounts of prior knowledge.
The structure of the rest of the paper is as follows. Section 2 reviews related work. Section 3 introduces the problem formulation and the federated graph learning framework. Section 4 presents the proposed gradient inversion attack algorithms, GLG, and Section 5 provides the corresponding theoretical analysis. Section 6 presents the experimental results. Section 7 concludes the paper and discusses future work.
2 Related Work
Gradient inversion attacks in FL. In terms of privacy leakage, communicating gradients throughout the training process in FL can reveal sensitive information about the participants. Deep leakage from gradients (DLG)(Zhu et al., 2019) demonstrated that sharing the gradients can leak private training data, including images and text. Specifically, Zhao et al. (2020a) proposes that the ground truth labels can be directly inferred from the gradients. Consequently, ensuing works improve the attack techniques on large batches of user data (Yin et al., 2021; Fowl et al., 2022b). A few recent works study reconstructing text data in federated learning. To facilitate the text reconstruction, Boenisch et al. (2021); Fowl et al. (2022a) reconstruct text data using a strong threat model in which the server is malicious and is able to manipulate the training model’s weights. In contrast, Gupta et al. (2022) studies the feasibility of reconstructing text from large batch sizes by leveraging the memorization capability of Language Models during federated learning. Recent work Petrov et al. (2024) introduces DAGER, the first exact gradient inversion attack for transformers by leveraging the low-rankness of self-attention layer gradients. The majority of the works are based on either image or text data.
Gradient inversion attacks in FGL. Reconstructing the data, including nodal features and graph structure, in FGL remains underexplored compared to the gradient inversion attacks in image or text domains. Since the input of GNN includes nodal features and graph structure, directly applying DLG or its enhanced methods in the graph domain is either infeasible or leads to significant performance degradation. Particularly in GCN, DLG consistently struggles to reconstruct the graph structure, which is essential for graph data. The only concurrent work, Drencheva et al. (2025), proposes GRAIN that introduces a gradient inversion attack tailored to GCN (Kipf & Welling, 2016) and GAT (Veličković et al., 2018). GRAIN employs a depth-first search strategy to identify subgraphs within gradients, iteratively filtering out unlikely candidates through a span check. By leveraging degree information in the features, GRAIN generates its final prediction. However, nodal degree information is not inherently accessible to attackers during GNN training. In the absence of degree information, neither theoretical nor experimental guarantees of GRAIN can keep reliable. Given the limited exploration of gradient inversion attacks in FGL, we propose a method, GLG, that reconstructs both nodal features and graph structure with fewer constraints. GLG does not rely on explicit degree information, making it more generalizable. To facilitate comparison, we summarize the key differences between DLG, GRAIN, and our method in Table 1. The detailed comparison is provided in Appendix H.
Table 1: Comparison of DLG, GRAIN and GLG. ✓and ✗indicate whether the method can reconstruct the corresponding graph data. Prior knowledge refers to any other information required besides the gradients.
Method Nodal feature Graph structure Prior knowledge
DLG None
GRAIN Node degree
GLG None
@@ 3.1 GNN Frameworks
We first introduce GraphSAGE and Graph Convolutional Network (GCN) frameworks for both node and graph classification tasks below.
@@ 4 Gradient Inversion Attack for GNNs
We consider the scenario where there is an honest-but-curious server, i.e., the server has access to the model gradient with respect to the client’s private graph data. Different attack scenarios are next discussed, where the server may also have access to different amounts of information.
@@ 4.1 Threat Model
The attacker is an honest-but-curious server that aims to reconstruct the user’s graph data given the gradients of a GNN model. In addition to possessing knowledge of the gradients, the attacker may also have access to varying amounts of information about the user’s private graph data. Hence we categorize the attacks based on the task as follows:
Node Attacker 1: has access to the gradients of a node classification model for the target node and reconstructs the nodal features of the target node and its neighbors.
Node Attacker 2: has access to the gradients of a node classification model for all the nodes in an egonet or subgraph. Specifically, the gradients are obtained for the loss function $\mathcal{L}_{v}\in\mathbb{R}^{1\times N}$ with respect to the neural networks weights, where each element $\mathcal{L}_{v_{i}}$ denotes the loss for node $v_{i}$. Unlike traditional classification tasks, the adjacency information is also significant for graph data. In this setting, the graph structure could be unknown. The honest-but-curious server would also try to reconstruct the nodal features and/or the graph structure. Based on the input, we categorize the attacker furhter as follows
  • Node Attacker 2-G: has access to the gradients of all the nodes in an egonet or subgraph along with the nodal features and reconstructs the underlying graph structure.
  • Node Attacker 2-N: has access to the gradients of all the nodes in an egonet or subgraph along with the graph structure and reconstructs the nodal features.
  • Node Attacker 2-GN: has access only to the gradients of all nodes in an egonet or subgraph and reconstructs the graph structure as well as the nodal features.
Graph Attacker: has access to the gradients of the graph for a graph classification model. Similar to Node Attacker 2, in this setting the honest-but-curious server would try to reconstruct the nodal features and/or the graph structure. Hence, we categorize the attack in more detail as follows
  • Graph Attacker-G: has access to the gradients along with the nodal features and reconstructs the underlying graph structure.
  • Graph Attacker-N: has access to the gradients along with the graph structure and reconstructs the nodal features.
  • Graph Attacker-GN: has access to the gradients and reconstructs both the nodal features and the graph structure.
Table 2: Overview of Threat Models: distinct attackers across seven settings. $(\mathbf{X})_{target}$ refers to nodal features of the target node and its neighbors. $\nabla_{\mathbf{W}}$ denotes the gradients of nodes, $\mathbf{X}$ represents features of all nodes in the subgraph, and $\mathbf{A}$ corresponds to the underlying graph structure.
Attacker setting Training task Accessible information Reconstructed data
Node Attacker 1 Node classification $\nabla_{\mathbf{W}}$ of the target node $(\mathbf{X})_{target}$
Node Attacker 2 G Node classification $\nabla_{\mathbf{W}}$ and $\mathbf{X}$ $\mathbf{A}$
Node Attacker 2 N Node classification $\nabla_{\mathbf{W}}$ and $\mathbf{A}$ $\mathbf{X}$
Node Attacker 2 GN Node classification $\nabla_{\mathbf{W}}$ $\mathbf{X}$ and $\mathbf{A}$
Graph Attacker G Graph classification $\nabla_{\mathbf{W}}$ and $\mathbf{X}$ $\mathbf{A}$
Graph Attacker N Graph classification $\nabla_{\mathbf{W}}$ and $\mathbf{A}$ $\mathbf{X}$
Graph Attacker GN Graph classification $\nabla_{\mathbf{W}}$ $\mathbf{X}$ and $\mathbf{A}$
@@ 4.2 Attack Mechanisms
In this section, GLG is introduced to attack a GNN framework and steal private user data using the gradients in a federated graph learning setting. Specifically, given the gradients of the model from a client for the $i$-th data sample $\nabla_{\mathbf{W}}\mathcal{L}_{c,i}$, the goal is to reconstruct the input graph data sample that was fed to the model. For simplicity, we denote $\nabla_{\mathbf{W}}\mathcal{L}_{c,i}$ as $\nabla_{\mathbf{W}}\mathcal{L}$ for the graph classification task, and as $\nabla_{\mathbf{W}}\mathcal{L}_{v}$ for the node classification task. In Zhu et al. (2019), the input data to a model is reconstructed by matching the dummy gradients with the gradients. The dummy gradients are obtained by feeding dummy input data to the model. Given the gradients at a certain time step as $\nabla_{\mathbf{W}}\mathcal{L}$, the input data is reconstructed by minimizing the following objective function
@@
In this work, it is assumed that the input label is known since in a classification task with cross-entropy loss, the labels can be readily inferred from the gradients (Zhao et al., 2020a). For further details regarding this please refer to Appendix B.
Feature Smoothness.
As shown in equations 2 and 4 Graph Neural Networks (GNNs) typically employ a feature aggregation mechanism wherein node representations are constructed by integrating features from neighboring nodes prior to weight multiplication at each network layer. However, the objective function proposed in equation 12 may be insufficient for simultaneously reconstructing both nodal features and graph structure. To address this limitation, we leverage intrinsic properties characteristic of real-world graphs that can enhance graph data reconstruction. Many real-world graph structures, such as social networks, exhibit a fundamental property of feature smoothness, wherein connected nodes tend to possess similar characteristics. To formally enforce and quantify this feature smoothness in the reconstructed graph data, we propose the following loss function, which is adapted from the approach introduced by Zhang et al. (2022):
$\mathcal{L}s$ $=tr(\mathbf{X}\mathbf{L}\mathbf{X}^{\top})$
$=\sum{i,j\in\mathcal{E}}(\frac{\mathbf{x}_{v_{i}}}{\sqrt{d_{i}}}-\frac{\mathbf{x}_{v_{j}}}{\sqrt{d_{j}}})$ (13)
The loss function captures the similarity between adjacent nodes by minimizing the normalized feature differences across connected nodes.
@@
where $\mathbb{D}$ is defined in equation 12, $\alpha$ and $\beta$ are hyperparameters. The iterative algorithms for reconstructing the private data from the gradients in a node classification task and a graph classification task are summarized in algorithms 1, 2, and 3, in Appendix A, respectively. Specifically, Node Attacker 1 utilizes algorithm 1 for reconstructing the nodal features of the target node and its neighbors. Node Attackers 2 (G, N and GN) utilizes algorithm 2 for reconstructing the nodal features and graph structure given the gradients of the loss function for each node in the egonet/subgraph. By default, algorithm 2 assumes that both nodal features and graph structure are unknown and tries to reconstruct both. However, only the unknown parameter will be optimized when either of the two is known. For instance, Node Attacker 2-G only optimizes for $\mathbf{A}$ in Line 9 of algorithm 2 while skipping Line 8 since the nodal features $\mathbf{X}$ is known. The same holds for Node Attacker 2-N. However, unlike the other two Node Attacker 2-GN optimizes for both. Similar logic holds for Graph Attacker (a, b and c) which utilizes algorithm 3. Note that algorithms 2 and 3 might seem similar since they are reconstructing the same variables. However, a key difference between the two is that algorithm 2 is attacking a node classification model while algorithm 3 attacks a graph classification model. In algorithm 2, the gradients of the loss function are computed individually for each node. This is because each node within the subgraph has an associated label. In contrast, algorithm 3 calculates gradients for the entire graph, as the graph has a single label.
Reconstructing the Adjacency Matrix: In all the scenarios where the attacker is trying to reconstruct the adjacency matrix, projected gradient descent is applied. Specifically, the gradient descent step in Line 9 of algorithms 2 and 3 entails a projection step, where each entry $(i,j)$ of the adjacency matrix is updated through the entry-wise projection operator defined as
@@
Finally, to reconstruct the binary adjacency matrix, we consider each entry $\hat{\mathbf{A}}_{ij}$ as the probability of any edge between nodes $v_{i}$ and $v_{j}$ and the reconstructed binary adjacency matrix is obtained by sampling from a Bernoulli distribution with the corresponding probability at the last iteration, see also Line 11 of algorithm 2 and algorithm 3.
Since the reconstructed adjacency matrix $\hat{\mathbf{A}}$ is obtained via sampling through Bernoulli distribution, the Frobenius regularizer has the effect of reducing the magnitude of entries of $\hat{\mathbf{A}}$. Consequently, this regularizer also contributes to the promotion of sparsity in the ultimately reconstructed adjacency matrix. This sparsity property is particularly relevant since real-world social networks are sparse.